What is the measure of one angle of a regular convex 100-gon?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

What is the measure of one angle of a regular convex 100-gon?
----
The sum of the exterior angles is 360
Each exterio is 360/100 = 3.6 degrees
---
Therefore, each interior angle is 180-3.6 = 176.4 degrees
========================
Cheers,
Stan H.
===============

Answer by Alan3354(69443)

What is the measure of one angle of a regular convex 100-gon?
-------
If it's regular, it has to be convex.
ext angles = 360/n = 3.6 degrees